OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 10x+5y=52.5 30x-7y=80.5

OpenStudy (anonymous):

multiply the first by -3, add to the second

OpenStudy (anonymous):

ok so 10x*-3

OpenStudy (anonymous):

ok now is the second part or the whole thing

OpenStudy (anonymous):

ok i messed up yet again so let me start afresh

OpenStudy (anonymous):

ok great

OpenStudy (anonymous):

here we go ready?

OpenStudy (anonymous):

\[10x+5y=52.5\] \[30x-7y=80.5\] multiply top by -3 get \[-30x-15y=-157.5\] \[30x-7y=80.5\]

OpenStudy (anonymous):

add the two equations together. the x's add up to zero and you get one equation in y, namely \[-22y=-77\] solve for y and get \[y=\frac{77}{22}=\frac{7}{2}=3.5\]

OpenStudy (anonymous):

now that we know y, replace y by 3.5 in either equation and solve for x

OpenStudy (anonymous):

\[10x+5y=52.5\] \[10x+5\times 3.5=52.5\] \[10x+17.5=52.5\] \[10x=35\] \[x=3.5\]

OpenStudy (anonymous):

so both x and y are 3.5 it doesn't usually work out that way

OpenStudy (anonymous):

ok got it now